What is the molarity of a 1.5 L solution containing 80.02g of sodium hydroxide.


1
Expert's answer
2021-04-09T06:26:39-0400

Solution:

sodium hydroxide - NaOH

Moles of NaOH = Mass of NaOH / Molar mass of NaOH

The molar mass of NaOH is 39.997 g mol-1.

Hence,

Moles of NaOH = (80.02 g) / (39.997 g mol-1) = 2.00 mol


Molarity of NaOH = Moles of NaOH / Liters of solution

Hence,

Molarity of NaOH = (2.00 mol) / (1.5 L) = 1.33 mol/L = 1.33 M

Molarity of NaOH = 1.33 M


OR (short form solution):

(80.02 g NaOH) × (1 mol NaOH / 39.997 g NaOH) × (1 / 1.5 L) = 1.33 M NaOH


Answer: The molarity of NaOH solution is 1.33 M.